英語 (English)

発信型英語 (Communicative English) (総科 ((総理 a)2 年))

岩佐憲明 · 非常勤講師 / 全学共通教育センター 2 単位後期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】This course introduces students to the wonders of philosophy by posing questions of values, knowledge, and reality. The course develops students’ philosophical thinking and improves their ability to express their thoughts in oral and written English.

【授業の概要】Each class briefly explains a philosophical question and a philosopher’s answer to it. Then, students are asked to think about the question philosophically, discuss it, and write their views in English.

【キーワード】 philosophy，哲学【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. To understand philosophical concepts and issues.

2. To develop philosophical thinking.

3. To improve the ability to express thoughts in oral and written English. 【授業の計画】

1. Introduction

2. Plato—Are you a fair and just person?

3. Mill—Will having fun make you happier than studying?

4. Kant—Should you ever tell a lie?

5. Heidegger—Do we control technology or does technology control us?

6. Descartes—Can you doubt that you exist?

7. Berkeley—Does a tree make a sound if it falls in a forest with no one around?

8. Hume—Are you certain that the law of gravity is really a law?

9. Wittgenstein—Can another person understand your feelings?

10. Dennett—Can computers think?

11. Parmenides—Can you think about nothing at all?

12. Democritus—Does anything ever happen by chance?

13. Aristotle—Are numbers and people equally real?

14. Locke—Are you the same person you were five years ago?

15. Hegel—Does anything depend on everything?

16. Review 【教科書】 PDF texts and MP3 audio are downloadable from https://spideroak.com (Click SHARE LOGIN, enter Share ID and Room Key, and click OPEN.)

【参考書等】 To be announced.

【成績評価の方法】 Grades are based on in-class writings and participation in discussion. 【再試験の有無】 n/a 【受講者へのメッセージ】 Read the text and listen to the audio before class. Bring the text,

writing paper, and a Japanese-English dictionary to class.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221208 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 岩佐 (niwasa@uchicago.edu) ���� (オフィスアワー: by appointment)

English

Communicative English (総科 ((総理 a)2 年))

Noriaki Iwasa · Part-time Lecturer / Center for General Education 2 units 後期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This course introduces students to the wonders of philosophy by posing questions of values, knowledge, and reality. The course develops students’ philosophical thinking and improves their ability to express their thoughts in oral and written English.

Outline⟩ Each class briefly explains a philosophical question and a philosopher’s answer to it. Then, students are asked to think about the question philosophically, discuss it, and write their views in English.

�������⟩ philosophy, 哲学 Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. To understand philosophical concepts and issues.

2. To develop philosophical thinking.

3. To improve the ability to express thoughts in oral and written English.

Schedule⟩

1. Introduction

2. Plato—Are you a fair and just person?

3. Mill—Will having fun make you happier than studying?

4. Kant—Should you ever tell a lie?

5. Heidegger—Do we control technology or does technology control us?

6. Descartes—Can you doubt that you exist?

7. Berkeley—Does a tree make a sound if it falls in a forest with no one around?

8. Hume—Are you certain that the law of gravity is really a law?

9. Wittgenstein—Can another person understand your feelings?

10. Dennett—Can computers think?

11. Parmenides—Can you think about nothing at all?

12. Democritus—Does anything ever happen by chance?

13. Aristotle—Are numbers and people equally real?

14. Locke—Are you the same person you were five years ago?

15. Hegel—Does anything depend on everything?

16. Review Textbook⟩ PDF texts and MP3 audio are downloadable from https://spideroak. com (Click SHARE LOGIN, enter Share ID and Room Key, and click OPEN.)

Reference⟩ To be announced.

Evaluation Criteria⟩ Grades are based on in-class writings and participation in

discussion. Re-evaluation⟩ n/a Message⟩ Read the text and listen to the audio before class. Bring the text,

writing paper, and a Japanese-English dictionary to class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221208 Contact (Office-Hour, Room, E-mail)⟩

⇒ Iwasa (niwasa@uchicago.edu) ���� (Office Hour: by appointment)